Write the standard form of the equation of the line parallel to the graph of 9x+3y=6 that passes through (0,3).
Answer by ikleyn(52787)

.

Any straight line in a coordinate plane, parallel to the given line has a standard form

    9x + 3y = d,      (*)

where "c" is a constant.



To find the parallel line, which goes through the point (0,3), substitute its coordinates x= 0 and y= 3

into equation (*).



You will get then

    c = 9*0 + 3*3 = 9.



So, a standard form equation under the problem's question is

    9x + 3y = 9,


or any other equivalent equation of this form, for example, this one

    3x + y = 3.      ANSWER

Solved.